Calculate solubility of AgI in 0.1 M NH3 ?


1
Expert's answer
2020-10-12T13:35:12-0400

Ks (AgI) = 8,3 ∙ 10-17

AgI = > Ag+ + I- Ksp = [Ag+ ] [ I-] = 8.3 * 10 ^ - 17

Ag+ + 2NH3 <=> Ag(NH3)2+

Kf = [Ag(NH3)2+] / [Ag+] [ NH3-] ^2 = 1.7 * 10 ^ 7

this is complexation

one Ag + that dissolves one I - ion is released

I- = [Ag(NH3)2+] = c

solving for [Ag+]

c^2 = Ksp * Kf * [NH3] ^2

c^2 = 8.3 * 10 ^-17 * 1.7*10 ^7 * (0.1-2c) ^2

c = -5.644 * 10 ^ -10 moles per liter